# Visualizing Quantum States: A Pilot Study on Problem Solving in Quantum Information Science Education

### v2

@@ Abstract
In the rapidly evolving interdisciplinary field of quantum information science and technology, a big obstacle is the need to understand high-level mathematics to solve complex problems. Current findings in educational research suggest that incorporating visualizations in settings of problem solving can have beneficial effects on students’ performance and cognitive load compared to solely relying on symbolic problem solving content. Visualizations like the (dimensional) circle notation enable us to represent not only single-qubit but also complex multi-qubit states, entanglement, and quantum algorithms. In this pilot study, we aim to take a first step to identify in which contexts students benefit from the presentation of visualizations of single- and multi-qubit systems in addition to mathematical formalism. For this purpose, we propose test items and a complete methodology to assess students’ performance and cognitive load when solving problems. This is a pilot investigation with a large breadth of questions intended to generate hypotheses and guide larger-scale but more concrete studies in the future. Specifically, we compare two approaches: using the mathematical-symbolic Dirac Notation (DN) alone and using it accompanied by the (dimensional) circle notation. Analyzing time and performance, we find that most of the test items are appropriate for a heterogeneous target group as they can differentiate between the participants in regard to performance and time taken. In general, the A-B crossover structure of the study is suitable for investigating the benefits of visualization for problem solving with respect to length and feasibility. Future studies should, however, be more restricted in context because of the observable student and context dependence, with special interest lying in the further investigation of the Hadamard gate, the CNOT gate, and entanglement in multi-qubit systems.
@@ I Introduction
The interdisciplinary nature of the field of Quantum Information Science and Technology ( Quantum Information Science & Technology (QIST)) necessitates endeavors to educate workforce with many different backgrounds [1]. This implies that the quantum education community needs to find ways to introduce the field of QIST to a broad audience. One way of facilitating such an introduction is the use of suitable representations, including visualization, to convey non-trivial quantum concepts.
@@
Visualization is generally considered a useful tool in Science, Technology, Engineering & Mathematics (STEM) education [2, 3, 4, 5, 6, 7] and in QIST education in particular [8, 9, 10]. The field of QIST is based on the utilization of systems of multiple qubits, but the effectiveness of presenting visualizations of qubit states in problem solving, in terms of performance [11] and cognitive load [12], is currently unknown. Knowing under which conditions and in which contexts learners benefit from the depiction of visualizations, we would be able to offer QIST educators valuable advice for teaching by being able to recommend the most effective ways of utilizing visualizations. This is an important step towards solving the challenge of high heterogeneity of target groups in QIST education. It is unknown whether and under which conditions learning and problem solving are supported by providing visualization in addition to the mathematical symbolic representation (e.g., the DN) in QIST education. We identify three relevant questions in the context of visualizations in QIST education.
1.
In what contexts does incorporating visualizations support students’ performance in task solving on quantum operations and entanglement?
2.
How does the inclusion of visualizations affect cognitive load?
3.
How are performance and cognitive load affected by representational competence?
Currently, there are no validated multiple choice test instruments designed to answer these questions. In this study, we propose test instruments and a novel study design to answer these questions. We also present preliminary results, discussing the adequacy of the questions and conclude with important lessons for future studies. In this way, we take a first step in the investigation of whether visualizations provide a benefit to solve problems in QIST when provided in addition to the DN and under what circumstances such benefits occur. Further studies can build on test instruments, survey design, and preliminary results to answer the research questions above. In this way, the primary goal of this work is to generate hypotheses and identify promising avenues for further investigation, thereby laying the groundwork for more focused and robust studies on the effective integration of visualizations in quantum education.

### v3

@@ Abstract
In the rapidly evolving interdisciplinary field of quantum information science and technology, a major obstacle is the need to understand advanced mathematics to solve complex problems. Current findings in educational research suggest that incorporating visualizations into problem-solving settings can have beneficial effects on students’ performance and cognitive load compared to relying solely on symbolic problem-solving content. Visualizations like the (dimensional) circle notation enable us to represent not only single-qubit but also more complex multi-qubit states, entanglement, and quantum algorithms. In this pilot study, we aim to take an initial step toward identifying the contexts in which students benefit from the presentation of visualizations of single- and multi-qubit systems in addition to mathematical formalism. For this purpose, we propose a set of test items and a comprehensive methodology to assess students’ performance and cognitive load when solving problems. This is a pilot investigation with a large breadth of questions intended to generate hypotheses and guide larger-scale, more focused studies in the future. Specifically, we compare two approaches: using the mathematical-symbolic Dirac Notation alone and using it in combination with the (dimensional) circle notation. In surveys in one-, two- and three-qubit systems, we gather qualitative data from five, five and two think-aloud interviews, identifying problems that students encounter and their problem-solving strategies. In addition, we analyze quantitative data (performance and cognitive load) from 23, 27 and 17 participants in surveys on one-, two- and three-qubit systems recruited mainly from our quantum computing lectures. We find that most of the test items are appropriate for a heterogeneous target group, as they can differentiate between participants in terms of performance and time taken. In general, the A-B crossover structure of the study is suitable for investigating the benefits of visualization for problem solving with regard to length and feasibility. Future studies should, however, be narrower in scope, given the observable dependence on student characteristics and context, with particular interest lying in the further investigation of the Hadamard gate, the CNOT gate, and Entanglement in multi-qubit systems.
@@ I Introduction
The interdisciplinary nature of the field of Quantum Information Science and Technology (QIST) necessitates concerted efforts to educate a workforce with diverse academic backgrounds [1]. This implies that the quantum education community needs to find ways to introduce the field of QIST to a broad audience. As a foundational concept, qubit systems are inherently interdisciplinary and span all of QIST. The “qubit-approach” can therefore be used to abstractly introduce a new, heterogeneous audience to the field. In consequence, the question of how best to make this approach as accessible as possible is raised. Fortunately, existing didactical theory provides a foundation from which to draw ideas.
@@
Visualization has been shown to be a useful tool in Science, Technology, Engineering & Mathematics (STEM) education [2, 3, 4, 5, 6, 7]. In QIST education in particular, visualization has been proposed to be useful for learning [8, 9]. The field of QIST is based on the utilization of systems of multiple qubits, but the effectiveness of presenting visualizations of qubit states in problem solving, in terms of performance [10] and cognitive load [11], is currently unknown. Understanding the contexts and conditions under which learners benefit from such visualizations would enable us to provide QIST educators with empirically grounded recommendations for how to most effectively integrate visualizations into teaching. This is an important step towards solving the challenge of high heterogeneity of target groups in QIST education. It remains unclear whether, and under which conditions, learning and problem solving are supported by providing visualizations alongside the mathematical–symbolic representation (e.g., the Dirac Notation (DN)) in QIST education. One theoretical construct that may offer explanatory insight is that of representational competence [12]. We identify three overarching questions in the context of visualizations in QIST education: In what contexts does incorporating visualizations support students’ performance in task solving on quantum operations and entanglement? How does the inclusion of visualizations affect cognitive load? How are performance and cognitive load affected by representational competence?
In order to answer these questions, appropriate and validated test instruments on task solving in QIST are needed. However, validated test instruments in the context of QIST education, and in particular on quantum operations and entanglement, are scarce in single-qubit systems and do, to our knowledge, not exist for multi-qubit systems. As an essential first step in the investigation of these questions, we conducted a pilot study validating relevant test instruments to investigate the overarching questions. In this way, we initiate an investigation into whether visualizations provide a benefit to solve problems in QIST when provided in addition to the DN and under what circumstances such benefits occur. The instruments, survey design, and preliminary results presented here lay the groundwork for future, more targeted studies that can build on this foundation.